# Counterfactual Sensitivity in Quantitative Spatial Models

### v2

@@ 3.4 Widely Applicable Default Approach
To provide a widely applicable default approach, consider the setting where we can write $T=g_{T}\left(\left\{F_{ij}\right\}\right)$ for a known function $g_{T}$ and a set of positive flows between locations, $\left\{F_{ij}\right\}$, a setup that is commonplace in quantitative spatial models (Costinot and Rodríguez-Clare, 2014; Redding and Rossi-Hansberg, 2017). Given the observed noisy flows $\left\{\tilde{F}_{ij}\right\}$, our goal is to find the posterior distribution $T|\left\{\tilde{F}_{ij}\right\}$. This requires a measurement error model and a prior on the true flows $\left\{F_{ij}\right\}$.
@@
In the default approach I recommend, both the prior on the true flows and the measurement errors are assumed to be log-normally distributed. This implies that posterior of the true flows given the noisy flows will also be log-normally distributed. This conjugacy is convenient, as it is computationally cheap to draw from the normal distribution. [^5] One could potentially enrich this by using a mixture of normals, in which case we will still have conjugacy. Furthermore, I assume that the prior mean exhibits a gravity relationship, for which there is strong empirical evidence, as for example outlined in Allen and Arkolakis (2018). This is summarized in the following assumption:
@@
It follows that the posterior for the true log flow between location $i$ and $j$, $\log F_{ij}$, given its noisy version, $\log\tilde{F}_{ij}$ is given by
@@
for $i,j=1,...,n$. Then, a default procedure for obtaining a credible set for $T$ is summarized in the following algorithm:
Algorithm 1 Credible set for $T=g_{T}\left(\left\{F_{ij}\right\}\right)$
@@
      Find estimates $\left\{\hat{\varsigma}^{2}\right\}$ using a specific data structure (as in Section 4) or domain knowledge (as in Section 5).
@@
      Obtain estimates $\left(\hat{\beta},\left\{\hat{\alpha}_{i}^{\mathrm{orig}}\right\},\left\{\hat{\alpha}_{i}^{\mathrm{dest}}\right\}\right)$ from the regression $\log\tilde{F}_{ij}=\beta\log\mathrm{dist}_{ij}+\alpha_{i}^{\mathrm{orig}}+\alpha_{j}^{\mathrm{dest}}+\phi_{ij}$, for $\phi_{ij}$ an error term.
@@
    Take $B$ draws (choose $B$ and $\gamma$ such that $\gamma/2\cdot B\in\mathbb{N}$) from the estimated posterior $\log F_{ij}|\log\tilde{F}_{ij}$ that uses $\hat{\vartheta}$ for $i,j=1,...,n$ and indicate them by $\log F_{ij,1},...,\log F_{ij,B}$ for $i,j=1,...,n$.
@@
    For $b=1,...,B$, compute $T_{b}=g_{T}\left(\left\{F_{ij,b}\right\}_{i,j=1}^{n}\right)$.
@@
    Sort these draws to obtain $\left\{T^{(b)}\right\}_{i=1}^{B}$ with $T^{(1)}\leq T^{(2)}\leq...\leq T^{(B)}$.
@@
    Report
    $\mathcal{C}_{T}^{\gamma}=\left[T^{(\gamma/2\cdot B)},T^{(\left(1-\gamma/2\right)\cdot B)}\right].$
@@
for $i,j=1,...,n$ and $t=1,...,\bar{t}$, which yields the posterior
@@
for $i,j=1,...,n$. This implies that now $\vartheta=\left(\left\{\beta_{t}\right\},\left\{\alpha_{i,t}^{\mathrm{orig}}\right\},\left\{\alpha_{i,t}^{\mathrm{dest}}\right\},\left\{s_{ij}^{2}\right\},\left\{\varsigma_{ij}^{2}\right\}\right)$. For $t=1,...,\bar{t}$, the estimates $\left(\hat{\beta}_{t},\left\{\hat{\alpha}_{i,t}^{\mathrm{orig}}\right\},\left\{\hat{\alpha}_{i,t}^{\mathrm{dest}}\right\}\right)$ are obtained from the regression $\log\tilde{F}_{ij,t}=\beta_{t}\log\mathrm{dist}_{ij,t}+\alpha_{i,t}^{\mathrm{orig}}+\alpha_{j,t}^{\mathrm{dest}}+\phi_{ij,t}$, and we have $\hat{s}_{ij}^{2}=\hat{r}_{ij}^{2}-\hat{\varsigma}_{ij}^{2}$, where $\hat{r}_{ij}^{2}$ denotes the sample variance of the residuals $\left\{\log\tilde{F}_{ij,t}-\hat{\beta}_{t}\log\mathrm{dist}_{ij}-\hat{\alpha}_{i,t}^{\mathrm{orig}}-\hat{\alpha}_{j,t}^{\mathrm{dest}}\right\}_{t=1}^{\bar{t}}$.
@@
One can verify how reasonable the normality assumption on the prior is by comparing the histogram of the normalized residuals
$\left\{\frac{\log\tilde{F}_{ij}-\left\{\hat{\beta}\log\mathrm{dist}_{ij}+\hat{\alpha}_{i}^{\mathrm{orig}}+\hat{\alpha}_{j}^{\mathrm{dest}}\right\}}{\hat{s}}\right\}$
with the probability density function of a standard normal. To check the reasonableness of the gravity prior, we can look at the adjusted R-squared of the gravity model and, following Allen and Arkolakis (2018), plot the log flows against the log distance, after partitioning out the origin and destination fixed effects. In Appendices D and E I perform both these checks for my applications.
@@
Specifying a measurement error model is difficult and one might be worried about misspecification. For the normal-normal model, we can use results on prior density-ratio class robustness developed in Geweke and Petrella (1998) to find worst-case bounds over a neighborhood that contains distributions that are not too far away from the assumed normal distribution for the measurement error model. The details can be found in Appendix F.
@@ 3.4.1 Running Example: Armington Model (Continued)
Consider again the Armington model where trade flows are noisily measured. This fits the setting from Section 3.4 and the prior and measurement error model are as in Equation (12). We can apply Algorithm 1 in which the estimated posterior of the true trade flow $\log F_{ij}$ given the noisy trade flow $\log\tilde{F}_{ij}$ is
@@
for $i,j=1,...,n$. Focusing again on the change in welfare of the first country which uses the estimated trade elasticity, $\hat{C}_{1}^{E}$, we can use the posterior in Equation (14) to find a credible set that covers $\left[L_{\hat{C}_{1}^{E}}^{\beta},U_{\hat{C}_{1}^{E}}^{\beta}\right]$ from Equation (8) with prescribed probability. Because we know that in turn $\hat{C}_{1}^{E}$ lies in $\left[L_{\hat{C}_{1}^{E}}^{\beta},U_{\hat{C}_{1}^{E}}^{\beta}\right]$ with prescribed probability, this allows us to then find a credible set for $\hat{C}_{1}^{E}$ as in Remark 1.
@@ 4.1 Model and Counterfactual Question of Interest
The empirical application of Adao, Costinot, and Donaldson (2017) investigates the effects of China joining the WTO, the so-called China shock. Specifically, the authors examine what would have happened to China’s welfare if China’s trade costs would stay constant at their 1995 levels. They consider $n$ countries and $\bar{t}$ time periods. Denoting with $\tau_{ij,t}$ the trade cost between country $i$ and country $j$ in period $t$, this implies the following counterfactual question of interest:
@@
Here, welfare is defined as the percentage change in income that the representative agent in China would be indifferent about accepting instead of the counterfactual change in trade costs from $\left\{\tau_{ij,t}\right\}$ to $\left\{\hat{\tau}_{ij,t}\tau_{ij,t}\right\}$. The details of the model can be found in Appendix D.1. The key insight is that we can express the estimated change in China’s welfare in period $t$, denoted by $\hat{W}_{\mathrm{China},t}^{E}$, as a function of all the bilateral trade flows in different periods $\left\{F_{ij,t}\right\}$ and the estimated trade elasticity. Since the estimated trade elasticity is itself a function of the bilateral trade flows, we can write
$\hat{W}_{\mathrm{China},t}^{E}=g_{\hat{W}_{\mathrm{China},t}}\left(\left\{F_{ij,t}\right\};\tilde{\varepsilon}\left(\left\{F_{ij,t}\right\}\right)\right),$ (15)
for known functions $\tilde{\varepsilon}:\mathbb{R}_{+}^{Tn\left(n-1\right)}\rightarrow\mathbb{R}$ and $g_{\hat{W}_{\mathrm{China},t}}:\mathbb{R}_{+}^{Tn\left(n-1\right)}\rightarrow\mathbb{R}$. We can then apply Algorithm 1 and Remark 1 to find a credible set for $\hat{W}_{\mathrm{China},t}^{E}$.
@@ 4.2 Measurement Error Model and Prior
The model fits the panel setting from Section 3.4 as outlined in Remark 2, and the prior and measurement error model are as in Equation (13). For estimation of $\vartheta$, I use the distance dataset $\left\{\mathrm{dist}_{ij}\right\}_{i\neq j}$ from Mayer and Zignago (2011) and the mirror trade dataset from Linsi, Burgoon, and Mügge (2023). This mirror trade dataset has two estimates of each bilateral trade flow, both as reported by the exporter and as by the importer. I interpret this as observing two independent noisy observations per time period for each bilateral trade flow: $\left\{\left\{\tilde{F}_{ij,t}^{1},\tilde{F}_{ij,t}^{2}\right\}_{t=1}^{T}\right\}_{i\neq j}$.
Then, to estimate $\varsigma_{ij}^{2}$, note that
@@
is an unbiased estimator of $\varsigma_{ij}^{2}$ for $i,j=1,...,n$. We can then estimate the rest of $\vartheta$ as outlined in Remark 2, adapting for the fact that we have two observations for each bilateral trade flow. For completeness, Appendix D.2 goes through the full procedure to obtain $\hat{\vartheta}=\left(\left\{\hat{\beta}_{t}\right\},\left\{\hat{\alpha}_{i,t}^{\mathrm{orig}}\right\},\left\{\hat{\alpha}_{i,t}^{\mathrm{dest}}\right\},\left\{\hat{s}_{ij}^{2}\right\},\left\{\hat{\varsigma}_{ij}^{2}\right\}\right)$.
To reduce the variability and leverage country information for $\left\{\hat{\varsigma}_{ij}^{2}\right\}$ and $\left\{\hat{s}_{ij}^{2}\right\}$, I fit the models:
@@
for $i,j=1,...,n$, with $\gamma_{i}^{\mathrm{exp}}$, $\gamma_{j}^{\mathrm{imp}}$, $\delta_{i}^{\mathrm{exp}}$ and $\delta_{i}^{\mathrm{imp}}$ country-exporter and country-importer fixed effects and $u_{ij}^{\varsigma}$ and $u_{ij}^{s}$ error terms. Then, rather than using $\hat{\varsigma}_{ij}^{2}$ and $\hat{s}_{ij}^{2}$ I will use the fitted values $\mathring{\varsigma}_{ij}^{2}=e^{\hat{\gamma}_{i}^{\mathrm{exp}}+\hat{\gamma}_{j}^{\mathrm{imp}}}$ and $\mathring{s}_{ij}^{2}=e^{\hat{\delta}_{i}^{\mathrm{exp}}+\hat{\delta}_{j}^{\mathrm{imp}}}$. Finally, the estimated posterior of the true trade flow $\log F_{ij,t}$ given the noisy trade flow $\log\tilde{F}_{ij,t}$ then is given, for $i,j=1,...,n$, by
$\mathcal{N}\left(\frac{\mathring{s}_{ij}^{2}}{\mathring{s}_{ij}^{2}+\mathring{\varsigma}_{ij}^{2}}\log\tilde{F}_{ij,t}+\frac{\mathring{\varsigma}_{ij}^{2}}{\mathring{s}_{ij}^{2}+\mathring{\varsigma}_{ij}^{2}}\left\{\hat{\beta}_{t}\mathrm{dist}_{ij}+\hat{\alpha}_{i,t}^{\mathrm{exp}}+\hat{\alpha}_{j,t}^{\mathrm{imp}}\right\},\left(\frac{1}{\mathring{s}_{ij}^{2}}+\frac{1}{\mathring{\varsigma}_{ij}^{2}}\right)^{-1}\right).$ (17)

### v3

@@ 3.4 Widely Applicable Default Approach for Uncertainty Quantification
Concretely, consider the setting where we can write $\hat{k}=g_{\hat{k}}\left(\left\{\log F_{ij}\right\};\theta\right)$, for $\left\{F_{ij}\right\}$ a set of positive flows between locations. We have an estimator $\tilde{\theta}\left(\left\{\log F_{ij}\right\}\right)$ with estimated sampling variance $\tilde{\Sigma}\left(\left\{\log F_{ij}\right\}\right)$. This setup is commonplace in quantitative trade and spatial models (Costinot and Rodríguez-Clare, 2014; Redding and Rossi-Hansberg, 2017; Proost and Thisse, 2019).
@@
Assume that both the prior distribution on the true flows and the measurement errors are log-normally distributed. This implies that the posterior distribution of the true flows given the noisy flows will also be log-normally distributed. This conjugacy is convenient, as it is computationally cheap to draw from the normal distribution. Furthermore, assume that the prior mean exhibits a gravity relationship, for which there is strong empirical evidence (Head and Mayer, 2014; Allen and Arkolakis, 2018). This is summarized in the following assumption:
@@
It follows that the posterior distribution for the true log flow between location $i$ and $j$, $\log F_{ij}$, given its noisy version, $\log\tilde{F}_{ij}$ is given by
@@
for $i,j=1,...,n$. The parameters $\vartheta=\left(\beta,\left\{\alpha_{i}^{\mathrm{orig}}\right\},\left\{\alpha_{i}^{\mathrm{dest}}\right\},s^{2},\varsigma^{2}\right)$ need to be estimated. One can find the estimator $\hat{\varsigma}^{2}$ using a specific data structure (as in Section 4) or domain knowledge (as in Section 5), and subsequently find the other estimators using Assumption 2. Having obtained $\hat{\vartheta}$, one can quantify uncertainty about $\hat{k}$ by finding the interval $\mathcal{C}^{1}$ as described in Section 3.3. Then, a default procedure for quantifying uncertainty about $\hat{k}$ is summarized in the following algorithm:
Algorithm 1 Uncertainty quantification about $\hat{k}=g_{\hat{k}}\left(\left\{\log F_{ij}\right\};\theta\right)$
@@
      Find the estimator $\hat{\varsigma}^{2}$ using a specific data structure (as in Section 4) or domain knowledge (as in Section 5).
@@
      Obtain estimates $\left(\hat{\beta},\left\{\hat{\alpha}_{i}^{\mathrm{orig}}\right\},\left\{\hat{\alpha}_{i}^{\mathrm{dest}}\right\}\right)$ from the regression $\log\tilde{F}_{ij}=\beta\log\mathrm{dist}_{ij}+\alpha_{i}^{\mathrm{orig}}+\alpha_{j}^{\mathrm{dest}}+\phi_{ij}$, with $\phi_{ij}$ an error term.
@@
    Take $B$ draws (choose $B$ and $\gamma$ such that $\gamma/2\cdot B\in\mathbb{N}$) from the estimated posterior distribution of $\log F_{ij}$ given $\log\tilde{F}_{ij}$ that uses $\hat{\vartheta}$ for $i,j=1,...,n$ and indicate them by $\log F_{ij,1},...,\log F_{ij,B}$ for $i,j=1,...,n$.
@@
    For $b=1,...,B$, sample $\theta_{b}$ from
    $\mathcal{N}\left(\tilde{\theta}\left(\left\{\log F_{ij,b}\right\}_{i,j=1}^{n}\right),\tilde{\Sigma}\left(\left\{\log F_{ij,b}\right\}_{i,j=1}^{n}\right)\right).$
@@
    For $b=1,...,B$, compute $\hat{k}_{b}=g_{\hat{k}}\left(\left\{\log F_{ij,b}\right\}_{i,j=1}^{n};\theta_{b}\right)$.
@@
    Sort these draws to obtain $\left\{\hat{k}^{(b)}\right\}_{b=1}^{B}$ with $\hat{k}^{(1)}\leq\hat{k}^{(2)}\leq...\leq\hat{k}^{(B)}$.
  6. 6.
    Report $\left[\hat{k}^{(\alpha/2\cdot B)},\hat{k}^{(\left(1-\alpha/2\right)\cdot B)}\right].$
@@
for $i,j=1,...,n$ and $t=1,...,T$. This implies the posterior distribution of $\log F_{ij,t}$ given $\log\tilde{F}_{ij,t}$ is given by
@@
for $i,j=1,...,n$. In this case we have $\vartheta=\left(\left\{\beta_{t}\right\},\left\{\alpha_{i,t}^{\mathrm{orig}}\right\},\left\{\alpha_{i,t}^{\mathrm{dest}}\right\},\left\{s_{ij}^{2}\right\},\left\{\varsigma_{ij}^{2}\right\}\right)$. For $t=1,...,T$, the estimates $\left(\hat{\beta}_{t},\left\{\hat{\alpha}_{i,t}^{\mathrm{orig}}\right\},\left\{\hat{\alpha}_{i,t}^{\mathrm{dest}}\right\}\right)$ are obtained from the regression $\log\tilde{F}_{ij,t}=\beta_{t}\log\mathrm{dist}_{ij,t}+\alpha_{i,t}^{\mathrm{orig}}+\alpha_{j,t}^{\mathrm{dest}}+\phi_{ij,t}$, and we have $\hat{s}_{ij}^{2}=\hat{r}_{ij}^{2}-\hat{\varsigma}_{ij}^{2}$, where $\hat{r}_{ij}^{2}$ denotes the sample variance of the residuals $\left\{\log\tilde{F}_{ij,t}-\hat{\beta}_{t}\log\mathrm{dist}_{ij}-\hat{\alpha}_{i,t}^{\mathrm{orig}}-\hat{\alpha}_{j,t}^{\mathrm{dest}}\right\}_{t=1}^{T}.$ [^7] In practice, $\hat{s}_{ij}^{2}$ can be negative. When this happens for only a very small share of flows, one can choose to omit these flows or set their prior variances equal to an arbitrarily small positive number. If this happens in the setting with a single prior variance as in Algorithm 1, I recommend to use a different prior or measurement error model.
@@
One can verify how reasonable the normality assumption on the prior and measurement error model is by comparing the histogram of the normalized residuals
$\left\{\left(\log\tilde{F}_{ij}-\left\{\hat{\beta}\log\mathrm{dist}_{ij}+\hat{\alpha}_{i}^{\mathrm{orig}}+\hat{\alpha}_{j}^{\mathrm{dest}}\right\}\right)/\hat{s}\right\}$
with the probability density function of a standard normal distribution. To further check the reasonableness of the gravity prior, we can look at the adjusted R-squared of the gravity model and, following Allen and Arkolakis (2018), plot the log flows against the log distance, after partitioning out the origin and destination fixed effects. In Appendices C and D I perform both these checks for my applications.
@@
Specifying a measurement error model is difficult and one might be worried about misspecification. For the normal-normal model, we can use results on prior density-ratio class robustness developed in Geweke and Petrella (1998) to find worst-case bounds over a neighborhood that contains distributions that are not too far away from the assumed normal distribution for the measurement error model. The details can be found in Appendix E.
@@ 3.4.1 Running Example: Armington Model (Continued)
Consider again the Armington model where bilateral trade flows are noisily measured. This fits the setting from Section 3.4 and the prior distribution and measurement error model are as in Equation (9). We can apply Algorithm 1, where the estimated posterior distribution of the true trade flow $\log F_{ij}$ given the noisy trade flow $\log\tilde{F}_{ij}$ is
@@
for $i,j=1,...,n$. Focusing again on the change in welfare of the first country which uses the estimated trade elasticity, we can use the posterior distribution in Equation (11) and Algorithm 1 to quantify uncertainty about $\hat{C}_{1}$
@@ 4.1 Model and Counterfactual Question of Interest
The empirical application of Adao, Costinot, and Donaldson (2017) investigates the effects of China joining the WTO, the so-called China shock. Specifically, the authors examine what would have happened to China’s welfare if China’s trade costs had stayed constant at their 1995 levels. They consider $n$ countries and $T$ time periods. Denoting with $\tau_{ij,t}$ the trade cost between country $i$ and country $j$ in period $t$, this implies the following counterfactual question of interest:
@@
Here, welfare is defined as the percentage change in income that the representative agent in China would be indifferent about accepting instead of the counterfactual change in trade costs from $\left\{\tau_{ij,t}\right\}$ to $\left\{\hat{\tau}_{ij,t}\tau_{ij,t}\right\}$. The details of the model can be found in Appendix C.1. The key insight is that we can express the change in China’s welfare in period $t$, denoted by $\hat{W}_{\mathrm{China},t}$, as a function of all the log bilateral trade flows in different periods $\left\{\log F_{ij,t}\right\}$ and the trade elasticity $\varepsilon$, which estimated by $\tilde{\varepsilon}\left(\left\{\log F_{ij,t}\right\}\right)$. Hence, we can write
$\hat{W}_{\mathrm{China},t}=g_{\hat{W}_{\mathrm{China},t}}\left(\left\{\log F_{ij,t}\right\};\varepsilon\right),$ (12)
for $t=1,...,T$ and for a known function $g_{\hat{W}_{\mathrm{China},t}}:\mathbb{R}_{+}^{Tn\left(n-1\right)}\rightarrow\mathbb{R}$. Then, conditional on a prior distribution for the true log bilateral flows $\left\{\log F_{ij,t}\right\}$ and a measurement error model, we can apply Algorithm 1 to quantify uncertainty about $\left\{\hat{W}_{\mathrm{China},t}\right\}$.
@@ 4.2 Measurement Error Model and Prior
The model fits the panel setting from Section 3.4 as outlined in Remark 1, and the prior distribution and measurement error model hence are as in Equation (10). For estimation of $\vartheta$, I use the distance dataset $\left\{\mathrm{dist}_{ij}\right\}_{i\neq j}$ from Mayer and Zignago (2011) and the mirror trade dataset from Linsi, Burgoon, and Mügge (2023). This mirror trade dataset has two estimates of each bilateral trade flow, both as reported by the exporter and as by the importer. I interpret this as observing two independent noisy observations per time period for each bilateral trade flow: $\left\{\left\{\tilde{F}_{ij,t}^{1},\tilde{F}_{ij,t}^{2}\right\}_{t=1}^{T}\right\}_{i\neq j}$. Then, to estimate $\varsigma_{ij}^{2}$, note that
@@
is an unbiased estimator of $\varsigma_{ij}^{2}$ for $i,j=1,...,n$. We can then estimate the other parameters as outlined in Remark 1, adapting for the fact that we have two observations for each bilateral trade flow. For completeness, Appendix C.2 goes through the full procedure to obtain $\hat{\vartheta}=\left(\left\{\hat{\beta}_{t}\right\},\left\{\hat{\alpha}_{i,t}^{\mathrm{orig}}\right\},\left\{\hat{\alpha}_{i,t}^{\mathrm{dest}}\right\},\left\{\hat{s}_{ij}^{2}\right\},\left\{\hat{\varsigma}_{ij}^{2}\right\}\right)$.
To reduce the variability and leverage country information for $\left\{\hat{\varsigma}_{ij}^{2}\right\}$ and $\left\{\hat{s}_{ij}^{2}\right\}$, I fit the models
@@
for $i,j=1,...,n$, with $\gamma_{i}^{\mathrm{exp}}$, $\gamma_{j}^{\mathrm{imp}}$, $\delta_{i}^{\mathrm{exp}}$ and $\delta_{i}^{\mathrm{imp}}$ country-exporter and country-importer fixed effects and $u_{ij}^{\varsigma}$ and $u_{ij}^{s}$ error terms. Then, rather than using $\hat{\varsigma}_{ij}^{2}$ and $\hat{s}_{ij}^{2}$ I will use the fitted values $\mathring{\varsigma}_{ij}^{2}=e^{\hat{\gamma}_{i}^{\mathrm{exp}}+\hat{\gamma}_{j}^{\mathrm{imp}}}$ and $\mathring{s}_{ij}^{2}=e^{\hat{\delta}_{i}^{\mathrm{exp}}+\hat{\delta}_{j}^{\mathrm{imp}}}$. Finally, the estimated posterior distribution of the true trade flow $\log F_{ij,t}$ given the noisy trade flow $\log\tilde{F}_{ij,t}$ then is given, for $i,j=1,...,n$, by
$\mathcal{N}\left(\frac{\mathring{s}_{ij}^{2}}{\mathring{s}_{ij}^{2}+\mathring{\varsigma}_{ij}^{2}}\log\tilde{F}_{ij,t}+\frac{\mathring{\varsigma}_{ij}^{2}}{\mathring{s}_{ij}^{2}+\mathring{\varsigma}_{ij}^{2}}\left\{\hat{\beta}_{t}\log\mathrm{dist}_{ij}+\hat{\alpha}_{i,t}^{\mathrm{exp}}+\hat{\alpha}_{j,t}^{\mathrm{imp}}\right\},\left(\frac{1}{\mathring{s}_{ij}^{2}}+\frac{1}{\mathring{\varsigma}_{ij}^{2}}\right)^{-1}\right).$ (14)